# Partial-immunity of two-photon correlation against wavefront distortion for spatially entangled photons

### v1

@@ I Introduction
High-dimensional entanglement is gaining popularity because of its ability to provide high bandwidth and error tolerance in quantum communication Cerf et al. (2002); Erhard, Krenn, and Zeilinger (2020). A commonly used method to generate high-dimensional entangled photons is spontaneous parametric down-conversion (SPDC) Klyshko (2018); Couteau (2018); Di Lorenzo Pires, Coppens, and van Exter (2011) in nonlinear crystals. Using SPDC, high-dimensional entanglement can be realized across various degrees of freedom, such as position-momentum entanglement Howell et al. (2004), energy-time entanglement Avenhaus et al. (2009), and angle-angular momentum entanglement Mair et al. (2001); Krenn et al. (2017). Of these, spatial entanglement is particularly appealing as spatial modes can be efficiently controlled using adaptive optics Boucher, Defienne, and Gigan (2021); Lib et al. (2024).
@@
When optical disorder is introduced in the path of high-dimensional spatially entangled photons, no interference is observed in the intensity distribution Peeters, Moerman, and Van Exter (2010). This occurs because the single-photon state comprises a statistical ensemble of multiple spatial modes and is, therefore, incoherent in nature Glauber (1963). Instead, the interference is observed in the coincidence space of two photons, which is referred to as two-photon interference Bouchard et al. (2020); Ghosh and Mandel (1987) and the random interference pattern as two-photon speckles Beenakker, Venderbos, and Van Exter (2009); Peeters, Moerman, and Van Exter (2010). The two-photon speckle indicates the randomization of correlation between the two photons. Hence, the scattering of photons poses a significant challenge in various quantum technologies, such as for instance, long-distance high-dimensional quantum key distribution Etcheverry et al. (2013); Otte et al. (2020); Zhou et al. (2019) and quantum imaging Unternährer et al. (2018); He et al. (2023), and the rectification of effects of disorder is a crucial field of research Valencia et al. (2020). To that end, earlier studies have shown that the two-photon interference caused by disorder in the near field of a nonlinear crystal can be controlled using a feedback mechanism Lib, Hasson, and Bromberg (2020) and a transmission matrix approach Courme et al. (2023). These interference patterns correlate with those produced by a coherent pump beam of half the wavelength when both entangled photons traverse the disorder Lib and Bromberg (2020). Additionally, the patterns correlate with the interference produced by a coherent beam of the same wavelength when only one of the two photons passes through the disorder Devaux et al. (2023). The two-photon correlations can be controlled in real-time using feedback from the coherent beam speckle Lib, Hasson, and Bromberg (2020). However, if the disorder is induced in the far field of the crystal, the two-photon wavefunction changes as it propagates to the far field Chan, Torres, and Eberly (2007). The two-photon interference pattern then does not correlate with the coherent beam’s interference pattern. In this case, restoring and controlling two-photon interference caused by disorder remains a challenge.
@@ II Theoretical framework
We introduce a phase modulation described by the transfer function $A_{d}(k)=\exp(i\phi(k))$ in the far field of the crystal plane. The wavefunction at the output of this phase modulation is given by:
@@
$\left.e^{-(k_{1}+k_{2})^{2}\sigma_{+}^{2}/{2}}e^{-(k_{1}-k_{2})^{2}\sigma_{-}^{2}/{2}}\langle 0|0\rangle\langle 0|0\rangle\right|^{2}$
@@
Furthermore, considering that the length scale of randomness in the phase modulation in k-space is much larger than $1/\sigma_{+}$ (i.e., weak disorder), the term $e^{-(k_{1}+k_{2})^{2}\sigma_{+}^{2}/2}$ can be effectively replaced by $\delta(k_{1}+k_{2})$ for $\sigma_{+}\gg\sigma_{-}$. Here, $\delta$ denotes the Dirac delta function, implying that the two photons have equal and opposite transverse momenta and thus appear at radially opposite points in the far-field plane. Under this approximation, the coincidence correlation function simplifies to:
@@
Figure 1: Numerical calculation of two-photon interference of spatially entangled photons and its comparison with interference pattern of auxiliary pump beam. The calculations are made using Eq. 6 and Eq. 7. (a)-(d) represent the amplitude of the initial Gaussian beam (red dashed curves) in k-space and the phase modulations (green solid curves), with the phase function mentioned in the respective plot. $\phi_{o}$ and $\phi_{e}$ in (c) and (d) are the odd and even components of the phase function in (b). (e)-(f) represent the two-photon interference pattern (red shade) in $(x_{1}-x_{2})/2$ and auxiliary pump intensity interference pattern (blue shade) in x for the same phase modulation. (a) & (e) correspond to the case when no phase is introduced, where the two interference patterns overlap exactly. (f) Single photon interference and two-photon correlations are affected by the phase. (g) For even phase, the effect on both is identical. Note also that the two-photon correlation patterns matches that in (f). (h) For odd phase, only the single photon interference is affected, while the two-photon correlation is preserved. The intensity and correlation patterns are normalised to maintain a constant area under the curve.
Any phase function $\phi(k)$ can be decomposed into even-parity ($\phi_{e}(k)$) and odd-parity ($\phi_{o}(k)$) components, i.e, $\phi_{e}(k)=(\phi(k)+\phi(-k))/2$ and $\phi_{o}(k)=(\phi(k)-\phi(-k))/2$. According to Eq. 5, only the even-parity component of the phase modulation contributes to the two-photon interference. The odd-parity component does not contribute to interference pattern because the equal and opposite phases of the two photons cancel each other out.
@@
To illustrate the theoretical results, we present a simple numerical example. Fig. 1 shows the calculation for two-photon interference using the phase modulation $\phi(k)=\cos(3k)+\sin(3k)$, and compares it with the interference pattern of the auxiliary pump beam. Fig. 1 (a)-(d) depict the amplitude of the beam in the k-space (red dashed curve, right Y-axis) and the phase introduced onto the beam (green solid curve, left Y-axis). The subplots illustrate (a) when no phase is introduced, (b) $\phi(k)=\cos(3k)+\sin(3k)$ is applied, (c) when only the even-parity component $\phi_{e}(k)$ is introduced, and (d) when only the odd-parity part $\phi_{o}(k)$ is introduced. Fig 1 (e)-(f) display the corresponding two-photon interference patterns (red shade) and auxiliary pump interference pattern (blue shade) for these four cases. These patterns show that the auxiliary pump beam is influenced by both the even-parity and odd-parity components, whereas the entangled photon pair experiences only the even-parity part. Specifically, when the phase is odd-parity (Fig. 1 (h)), the auxiliary pump exhibits an interference pattern, but the entangled photons show no two-photon interference. Conversely, when an even-parity phase modulation is introduced (Fig. 1 (g)), the interference pattern of the auxiliary pump matches the two-photon interference pattern in the variable $(x_{1}-x_{2})/2$. Also, the two-photon interference pattern corresponding to the total phase $\phi$ (Fig. 1 (f)) is identical to the interference pattern observed when only $\phi_{e}$ is applied (Fig. 1(g)). This confirms that only the even-parity component of the phase modulation contributes to the interference of entangled photons.
@@ III.1 Experimental Setup
Fig. 3 illustrates the experimental setup. A vertically polarized pump beam with a wavelength of $405~{}nm$ and a waist size of $\approx 425~{}\mu m$ is directed through a $5~{}mm$ long Type-0 periodically poled potassium titanyl phosphate (PPKTP) crystal. The crystal generates vertically polarized, degenerate photon pairs at a wavelength of $810~{}nm$ through SPDC. The emitted photon pairs exhibit spatial entanglement, with a Schmidt number of $\approx 700$ (see Appendix B). The far field of the crystal’s output face is imaged onto the DM using lens L2. The crystal is positioned on the front focal plane of L2, and the DM is located on the back focal plane. The focal length of L2 was carefully chosen to ensure the entire area of the DM is covered by photon pairs, maximizing its effective utilization. A polarizing beamsplitter (PBS) is used to separate the incident and reflected beams from the DM. Before L2, a half wave- plate (HWP) rotates the polarization of the $810~{}nm$ photons to horizontal, allowing them to be transmitted through the PBS and directed to the DM. A quarter wave-plate (QWP), placed between the PBS and DM, rotates the polarization back to vertical after reflection. A long-pass dichroic mirror (DMLP) separates the pump beam from the down-converted photons, directing them to separate detection paths. The pump beam is detected using a CCD camera, while the down-converted photons are captured using an electron-multiplying charge-coupled device (EMCCD). The deformations introduced by the DM are observed in the far field using lens L3, with the DM positioned at the front focal plane and the detectors at the back focal plane of L3. To block stray photons, an $810\pm 5~{}nm$ bandpass filter (not shown) is mounted in front of the EMCCD.
@@
Figure 3: Experimental setup. L: Lens, PPKTP: Periodically-poled potassium titanyl phosphate crystal, HWP: Half Wave-plate for $810~{}nm$, PBS: Polarizing beamsplitter, QWP: Quarter Wave-plate for $810~{}nm$, DM: Deformable mirror, DMLP: Longpass dichroic mirror, EMCCD: Electron multiplying charge coupled device, CCD: Charge coupled device. Lens L1 is inserted before the crystal to observe $405~{}nm$ interference pattern using CCD.
@@
In the initial setup (Fig. 3), the pump beam is collimated and then focused onto the deformable mirror (DM) by lens L2. Since the pump beam is focused onto the center of the DM, it does not undergo any deformations from the DM, and thus no interference pattern is observed. To capture the effects of the DM using the pump beam, we introduce the auxiliary pump by adding a lens combination with effective focal length $3~{}cm$ (represented here as L1) before the crystal. This focuses the beam onto the crystal. Lens L2 then collimates the auxiliary pump beam again and directs it to the DM. The term ’auxiliary pump’ is used here to indicate that we are no longer detecting the down-converted photons but instead analyzing the deformations of the pump beam itself, using the CCD camera. We verified that the disorder strength applied by the DM for the two wavelengths, $405~{}nm$ and $810~{}nm$, is comparable (see Appendix C). The two-photon interference pattern (without L1) from the entangled photons is recorded separately using the EMCCD. These measurements are taken consecutively, and the interference pattern of the auxiliary pump is compared with the two-photon interference pattern.
@@
The focal lengths of lenses L2, and L3 are $40~{}cm$, and $50~{}cm$, respectively. To ensure that both the auxiliary pump and the down-converted photons experience the same deformations, two conditions are met: (a) the beams should have the same size on the DM, achieved by selecting the appropriate focal length for L1, and (b) they must overlap perfectly on the DM plane, ensured by carefully aligning L1.
@@
Figure 4: Experimental results for no phase (first row) and two different configurations of odd-parity disorder (second and third row) implemented using DM. The left column shows the auxiliary pump interference pattern and the right column represents two-photon correlation in $x_{1}-x_{2}$ coincidence space corresponding to the same configuration. Odd parity disorder has no effect on two-photon correlation.
@@ III.2 Experimental Results
Figure 6: Result of numerical simulation with a two dimensional double Gaussian wavefunction in momentum variable illuminating a lossless disorder and observed in the far field. (a)-(d) represent auxiliary pump intensity, and (e)-(f) represents two-photon correlation in coincidence space as a function of difference in detector positions for the same disorder configuration. (a), (e) are the results without any disorder. (b), (f) are the results when a general phase is added to the wavefronts. (c), (g) are the results corresponding to the even-parity part of the general phase, and (d), (f) are the results for the odd-parity part. (i) and (f) show the 1D cross-section of the two-photon correlation corresponding to the above disorder configurations along $y_{1}-y_{2}=0$ and $y_{1}-y_{2}=0.64$ for comparison.
@@ IV Simulation
We perform numerical simulations by considering a two-dimensional double Gaussian wavefunction in momentum space and a Gaussian auxiliary pump illuminating a lossless disorder. Far-field interference patterns are observed for both by applying a Fourier transform to the field after the disorder, followed by taking the mode square. To observe the two-photon interference, the 4-dimensional wavefunction is projected along $x_{1}-x_{2}$ vs. $y_{1}-y_{2}$, revealing the spatial correlations between photon pairs. The Schmidt number of this simulated two-dimensional wavefunction is $\approx 1600$. Fig. 6 presents the results of numerical simulations. The first row of Fig. 6 shows the auxiliary pump interference patterns and the second row presents the two-photon interference patterns corresponding to the same random phase configurations. Fig. 6 (a) and  6 (e) correspond to the case with no disorder ($\phi(k)=0$). As expected, the auxiliary pump shows focusing in the far field, while the two photons show perfect correlation, which manifests as a sharp peak in the two-photon interference pattern. When a random phase disorder is introduced, the auxiliary pump field gets modified and exhibits a strong speckle pattern shown in Fig. 6 (b). Correspondingly, the correlation between the two photons is randomized, and a two-photon speckle pattern is observed in Fig. 6 (f). Given that the disorder affects the two systems differently, there is no visible pattern correlation between the two-photon speckle and the auxiliary pump speckle.

### v2

# Partial immunity of two-photon correlation against wavefront distortion for spatially entangled photons

@@ I Introduction
High-dimensional entanglement is gaining popularity because of its ability to provide high bandwidth and error tolerance in quantum communication Cerf et al. (2002); Erhard, Krenn, and Zeilinger (2020). A commonly used method to generate high-dimensional entangled photons is spontaneous parametric down-conversion (SPDC) Klyshko (2018); Couteau (2018); Di Lorenzo Pires, Coppens, and van Exter (2011) in nonlinear crystals. Using SPDC, high-dimensional entanglement can be realized across various degrees of freedom, such as position-momentum (spatial) entanglement Howell et al. (2004), energy-time entanglement Avenhaus et al. (2009), and angle-angular momentum entanglement Mair et al. (2001); Krenn et al. (2017). Of these, spatial entanglement is particularly appealing as spatial modes can be efficiently controlled using adaptive optics Boucher, Defienne, and Gigan (2021); Lib et al. (2024).
@@
When optical disorder is introduced in the path of high-dimensional spatially entangled photons, no interference is observed in the intensity distribution Peeters, Moerman, and Van Exter (2010). This occurs because the single-photon state comprises a statistical ensemble of multiple spatial modes and is, therefore, incoherent in nature Glauber (1963). Instead, the interference is observed in the coincidence space of two photons, which is referred to as two-photon interference Bouchard et al. (2020); Ghosh and Mandel (1987) and the random interference pattern as two-photon speckles Beenakker, Venderbos, and Van Exter (2009); Peeters, Moerman, and Van Exter (2010). The two-photon speckle indicates the randomization of correlation between the two photons. Hence, the scattering of photons poses a significant challenge in various quantum technologies, such as for instance, long-distance high-dimensional quantum key distribution Etcheverry et al. (2013); Otte et al. (2020); Zhou et al. (2019) and quantum imaging Unternährer et al. (2018); He et al. (2023), and the rectification of effects of disorder is a crucial field of research Valencia et al. (2020). To that end, earlier studies have shown that the two-photon interference caused by disorder in the near field of a nonlinear crystal can be controlled using a feedback mechanism Lib, Hasson, and Bromberg (2020) and a transmission matrix approach Courme et al. (2023). These interference patterns correlate with those produced by a coherent pump beam of half the wavelength when both entangled photons traverse the disorder Lib and Bromberg (2020). Additionally, the patterns correlate with the interference produced by a coherent beam of the same wavelength when only one of the two photons passes through the disorder Devaux et al. (2023). The two-photon correlations can be controlled in real-time using feedback from the coherent beam speckle Lib, Hasson, and Bromberg (2020). However, if the disorder is induced in the far field of the crystal, the two-photon wavefunction changes as it propagates to the far field Chan, Torres, and Eberly (2007). The two-photon interference pattern then does not correlate with the coherent beam’s interference pattern. In this case, restoring and controlling two-photon interference caused by disorder is more challenging Black et al. (2019).
@@ II Theoretical framework
We introduce a phase modulation, described by the transfer function $A_{d}(k)=\exp(i\phi(k))$, in the far field of the crystal plane. The wavefunction at the output of this phase modulation is given by:
@@
$\left.e^{-(k_{1}+k_{2})^{2}\sigma_{+}^{2}/{2}}e^{-(k_{1}-k_{2})^{2}\sigma_{-}^{2}/{2}}\right|^{2}\langle 0|0\rangle\langle 0|0\rangle$
@@
Furthermore, considering that the length scale of randomness in the phase modulation in k-space is much larger than $1/\sigma_{+}$ (i.e., weak disorder), the term $e^{-(k_{1}+k_{2})^{2}\sigma_{+}^{2}/2}$ can be effectively replaced by $\delta(k_{1}+k_{2})$ for $\sigma_{+}\gg\sigma_{-}$. Here, $\delta$ denotes the Dirac delta function, implying that the two photons have equal and opposite transverse momenta and thus appear at radially opposite points in the far field plane. Under this approximation, the coincidence correlation function simplifies to:
@@
Figure 1: Numerical calculations of two-photon interference of spatially entangled photons and its comparison with interference pattern of auxiliary pump beam. The calculations are made using Eq. 6 and Eq. 7. Panels (a)-(d) represent the amplitude of the initial Gaussian beam (red dashed curves) in k-space and the phase modulations (green solid curves), with the phase function mentioned in the respective plot. $\phi_{o}$ and $\phi_{e}$ in (c) and (d) are the odd and even components of the phase function in (b). Panels (e)-(h) represent the two-photon interference pattern (red shade) in $(x_{1}-x_{2})/2$ and auxiliary pump intensity interference pattern (blue shade) in x for the same phase modulation. (a) & (e) correspond to the case when no phase is introduced, where the two interference patterns overlap exactly. (f) Single photon interference and two-photon correlations are affected by the phase. (g) For even phase, the effect on both is identical. Note also that the two-photon correlation patterns matches that in (f). (h) For odd phase, only the single photon interference is affected, while the two-photon correlation is preserved. The intensity and correlation patterns are normalised to maintain a constant area under the curve.
The symmetry of the wavefuntion suggests that we should examine the odd-parity ($\phi_{o}(k)$) and even-parity ($\phi_{e}(k)$) components of $\phi(k)$ Black et al. (2019); Bonato et al. (2008), i.e, $\phi_{e}(k)=(\phi(k)+\phi(-k))/2$ and $\phi_{o}(k)=(\phi(k)-\phi(-k))/2$. According to Eq. 5, only the even-parity component of the phase modulation contributes to the two-photon interference. The odd-parity component does not contribute to interference pattern because the equal and opposite phases of the two photons cancel each other out.
@@
To illustrate the theoretical results, we present a simple numerical example. Fig. 1 shows the calculation for two-photon interference using the phase modulation $\phi(k)=\cos(3k)+\sin(3k)$, and compares it with the interference pattern of the auxiliary pump beam. Fig. 1 (a)-(d) depict the amplitude of the beam in the k-space (red dashed curve, right Y-axis) and the phase introduced onto the beam (green solid curve, left Y-axis). The subplots illustrate (a) when no phase is introduced, (b) $\phi(k)=\cos(3k)+\sin(3k)$ is applied, (c) when only the even-parity component $\phi_{e}(k)$ is introduced, and (d) when only the odd-parity part $\phi_{o}(k)$ is introduced. Fig 1 (e)-(h) display the corresponding two-photon interference patterns (red shade) and auxiliary pump interference pattern (blue shade) for these four cases. These patterns show that the auxiliary pump beam is influenced by both the even-parity and odd-parity components, whereas the entangled photon pair experiences only the even-parity part. Specifically, when the phase is odd-parity (Fig. 1 (h)), the auxiliary pump exhibits an interference pattern, but the entangled photons show no two-photon interference. Conversely, when an even-parity phase modulation is introduced (Fig. 1 (g)), the interference pattern of the auxiliary pump matches the two-photon interference pattern in the variable $(x_{1}-x_{2})/2$. Also, the two-photon interference pattern corresponding to the total phase $\phi$ (Fig. 1 (f)) is identical to the interference pattern observed when only $\phi_{e}$ is applied (Fig. 1(g)). This confirms that only the even-parity component of the phase modulation contributes to the interference of entangled photons.
@@ III.1 Experimental Setup
Fig. 3 illustrates the experimental setup. A vertically polarized pump beam with a wavelength of $405$ nm and a waist size of $\approx 425\mu$m is directed through a $5$ mm long Type-0 periodically poled potassium titanyl phosphate (PPKTP) crystal. The crystal generates vertically polarized, degenerate photon pairs at a wavelength of $810$ nm through SPDC. The emitted photon pairs exhibit spatial entanglement, with a Schmidt number of $\approx 700$ (see Appendix B). The far field of the crystal’s output face is imaged onto the DM using lens L2. The crystal is positioned on the front focal plane of L2, and the DM is located on the back focal plane. The focal length of L2 was carefully chosen to ensure the entire area of the DM is covered by photon pairs, maximizing its effective utilization. A polarizing beamsplitter (PBS) is used to separate the incident and reflected beams from the DM. Before L2, a half wave- plate (HWP) rotates the polarization of the $810$ nm photons to horizontal, allowing them to be transmitted through the PBS and directed to the DM. A quarter wave-plate (QWP), placed between the PBS and DM, rotates the polarization back to vertical after reflection. A long-pass dichroic mirror (DMLP) separates the pump beam from the down-converted photons, directing them to separate detection paths. The pump beam is detected using a CCD camera, while the down-converted photons are captured using an electron-multiplying charge-coupled device (EMCCD). The deformations introduced by the DM are observed in the far field using lens L3, with the DM positioned at the front focal plane and the detectors at the back focal plane of L3. To block stray photons, an $810\pm 5$ nm bandpass filter (not shown) is mounted in front of the EMCCD.
@@
Figure 3: Experimental setup. L: Lens, PPKTP: Periodically-poled potassium titanyl phosphate crystal, HWP: Half Wave-plate for $810$ nm, PBS: Polarizing beamsplitter, QWP: Quarter Wave-plate for $810$ nm, DM: Deformable mirror, DMLP: Longpass dichroic mirror, EMCCD: Electron multiplying charge coupled device, CCD: Charge coupled device. Lens L1 is inserted before the crystal to observe $405$ nm interference pattern using CCD.
@@
In the initial setup (Fig. 3), the pump beam is collimated and then focused onto the deformable mirror (DM) by lens L2. Since the pump beam is focused onto the center of the DM, it does not undergo any deformations from the DM, and thus no interference pattern is observed. To capture the effects of the DM using the pump beam, we introduce the auxiliary pump by adding a lens combination with effective focal length $3$ cm (represented here as L1) before the crystal. This focuses the beam onto the crystal. Lens L2 then collimates the auxiliary pump beam again and directs it to the DM. The term ’auxiliary pump’ is used here to indicate that we are no longer detecting the down-converted photons but instead analyzing the deformations of the pump beam itself, using the CCD camera. We verified that the disorder strength applied by the DM for the two wavelengths, $405$ nm and $810$ nm, is comparable (see Appendix C). The two-photon interference pattern (without L1) from the entangled photons is recorded separately using the EMCCD. These measurements are taken consecutively, and the interference pattern of the auxiliary pump is compared with the two-photon interference pattern.
@@
The focal lengths of lenses L2, and L3 are $40$ cm, and $50$ cm, respectively. To ensure that both the auxiliary pump and the down-converted photons experience the same deformations, two conditions are met: (a) the beams should have the same size on the DM, achieved by selecting the appropriate focal length for L1, and (b) they must overlap perfectly on the DM plane, ensured by carefully aligning L1.
@@
Figure 4: Experimental results for no phase (first row) and two different configurations of odd-parity disorder (second and third row) implemented using DM. Panels (a), (c) and (e) show the auxiliary pump interference pattern and panels (b), (d) and (f) represent two-photon correlation in $x_{1}-x_{2}$ coincidence space corresponding to the same configuration. Odd parity disorder has no effect on two-photon correlation.
@@ III.2 Experimental Results
Figure 6: Result of numerical simulation with a two-dimensional double Gaussian wavefunction in momentum variable illuminating a random phase mask and observed in the far field. (a)-(d) represent auxiliary pump intensity, and (e)-(h) represents two-photon correlation in coincidence space as a function of difference in detector positions for the same disorder configuration. (a), (e) are the results without any disorder. (b), (f) are the results when a general phase is added to the wavefronts. (c), (g) are the results corresponding to the even-parity part of the general phase, and (d), (h) are the results for the odd-parity part. (i) and (f) show the 1D cross-section of the two-photon correlation corresponding to the above disorder configurations along $y_{1}-y_{2}=0$ and $y_{1}-y_{2}=0.64$ for comparison.
@@ IV Simulation
We perform numerical simulations by considering a two-dimensional double Gaussian wavefunction in momentum space and a Gaussian auxiliary pump illuminating a random phase mask. Far field interference patterns are observed for both by applying a Fourier transform to the field after the disorder, followed by taking the squared modulus. To observe the two-photon interference, the 4-dimensional wavefunction is projected along $x_{1}-x_{2}$ vs. $y_{1}-y_{2}$, revealing the spatial correlations between photon pairs. The Schmidt number of this simulated two-dimensional wavefunction is $\approx 1600$. Fig. 6 presents the results of numerical simulations. The first row of Fig. 6 shows the auxiliary pump interference patterns and the second row presents the two-photon interference patterns corresponding to the same random phase configurations. Fig. 6 (a) and  6 (e) correspond to the case with no disorder ($\phi(k)=0$). As expected, the auxiliary pump shows focusing in the far field, while the two photons show perfect correlation, which manifests as a sharp peak in the two-photon interference pattern. When a random phase disorder is introduced, the auxiliary pump field gets modified and exhibits a strong speckle pattern shown in Fig. 6 (b). Correspondingly, the correlation between the two photons is randomized, and a two-photon speckle pattern is observed in Fig. 6 (f). Given that the disorder affects the two systems differently, there is no visible pattern correlation between the two-photon speckle and the auxiliary pump speckle.